Tom scored 75, 78, 87, 81 and 76 on five exams. What grade must Tom obtain on his sixth exam to have an average of 80?
Found 2 solutions by stanbon, nerdybill:
Answer by stanbon(75887)

Tom scored 75, 78, 87, 81 and 76 on five exams. What grade must Tom obtain on his sixth exam to have an average of 80?
-------------
Let the 6th grade be "x".
---
Equation:
Solve for "x".
(75+78+87+81+76+x)/6 = 80
======================
Cheers,
Stan H.

Answer by nerdybill(7384)

Let x = score on sixth exam
then
(75+78+87+81+76+x)/6 = 80
(75+78+87+81+76+x) = 480
(397+x) = 480
397+x = 480
x = 480-397
x = 83